Respected sir,
plz help me to slove this % problem.
1] At car delarship X,the total profil from sales increased by 10% over the previous year while the number of cars sold decreased by 10% over the previous year.Approximately what was the average percent increase in profit per car over the previous year?
waiting for yur reply.
Thanks.
Regards
Binita.

Answer by Theo(13342)

Your problem is:
1] At car delarship X,the total profit from sales increased by 10% over the previous year while the number of cars sold decreased by 10% over the previous year.
-----
Approximately what was the average percent increase in profit per car over the previous year?
-----
let P = total profit made the previous year.
let C = number of cars sold the previous year.
-----
Average Profit = Total Profit divided by number of cars sold.
-----
let A = average profit per car made the previous year.
Formula for calculating average profit per car from the previous year becomes:
A = P/C
-----
This year the total profit increased by 10% over the previous year.
This makes the total profit this year equal to P + .1*P = 1.1 * P
-----
This year the number of cars sold decreased by 10% over the previous year.
This makes the total number of cars sold this year equal to C - .1*C which equals .9 * C
-----
let B = average profit per car made this year.
Formula for calculating average profit per car this year becomes:
B = 1.1 * P / .9 * C
-----
This is the same as:
B = (1.1/.9) * (P/C)
-----
1.1/.9 = 1.22222222...
-----
This means that the average profit this year per car increased 22.222222... % over the average profit per car from the previous year.
-----
This is how that works.
A = average profit per car last year = P/C
B = average profit per car this year = (1.1*P)/(.9*C)
-----
(1.1P/.9C)/(P/C) is the same as:
(1.1P/.9C)*(C/P)
P/C cancels out C/P so you are left with 1.1/.9.
-----
The result is that B/A = 1.1/.9 = 1.22222222... which is the same as an increase of 22.22222... %
-----
B/A = 1.1/.9 becomes:
B = A*1.1/.9 when you solve for B by multiplying both sides of this equation by A.
-----
A*1.1/.9 becomes A*1.22222.... when you divide 1.1 by .9.
-----
A * 1.22222.... is the same as A + .22222... * A
-----
.22222... * A is the same as 22.22222% * A when you multiply .222 * 100% to turn it into a percent rather than a ratio.
-----
let me put this into numbers so you can see it better.
-----
let C = 1,000
let P = 5,000,000
-----
A = average profit per car = P/C = 5,000,000/1,000 = 5,000 dollars profit per car.
-----
This year the total profit went up 10% over last year.
P + 10% * P = 1.1 * P = 1.1 * 5,000,000 = 5,500,000
-----
this year the total number of cars sold went down 10% over last year.
C - 10% * C = .9 * C = .9 * 1,000 = 900
-----
B = average profit per car this year = 1.1*P/.9*C = 5,500,000/900 = 6111.11111... per car.
-----
6111.11111... / 5000 = 1.222222222...
This means that average profit per car this year increased by 22.22...% over last year.
-----
Rounding this to the nearest hundredth of a percent makes the answer equal to:
Average increase in profit per car this year equals 22.22%
-----